Find the first partial derivative of

f(x,y) = "(4x-y)\/(4x+y)" at the point (4,2) for both x and y.



1
Expert's answer
2022-05-13T03:03:14-0400

"f(x,y)=(4x-y)\/(4x+y)"

"f_x=\frac{(4x-y)_x'(4x+y)-(4x+y)_x'(4x-y)}{(4x+y)^2}=\frac{4(4x+y)-4(4x-y)}{(4x+y)^2}=\frac{8y}{(4x+y)^2}"

At the point (4,2)

"f_x=\frac{8(2)}{(4(4)+2)^2}=16/324=4/81"

"f_y=\frac{(4x-y)_y'(4x+y)-(4x+y)_y'(4x-y)}{(4x+y)^2}=\frac{-1(4x+y)-1(4x-y)}{(4x+y)^2}=\frac{-8x}{(4x+y)^2}"

At the point (4,2)

"f_y=\frac{-8(4)}{(4(4)+2)^2}=-32/324=-8/81"